An agent consumes quantity (x1,x2) of goods 1 and 2. Here is his utility function: ๐‘ˆ(๐‘ฅ1, ๐‘ฅ2) = 2 โˆ— 1 โˆ— ๐‘ฅ2 + 1, ๐‘1 = ๐‘2 = $1, โ„Ž๐‘–๐‘  ๐‘–๐‘›๐‘๐‘œ๐‘š๐‘’ ๐‘š = 20.

ย (a) Calculate the agentโ€™s Marshallian demand (x*1 , x*2 ).

(b) If the government put a $1 tax on x1, which increase p1 to $2, assume p2 and m do not change, what is the demand for x1?

(c) If the government collect tax on the agentโ€™s income, the amount of tax is the same as the tax revenue collected in (b), what is the agentโ€™s utility? Compare it with the agentโ€™s utility in (b).ย 


1
Expert's answer
2021-10-12T09:53:45-0400

Utility function isย "x_1^3x_2"ย with budget constraintย "p_1x_1+p_2x_2=m"

At optimum,ย ย "\frac{MU

X_1}{

MU

X_2}



=\frac{p_1}{

p_2}"

a) Therefore,

"MUx_1=3x_1^2x_2\\MUx_2=x_1^3\\At\space optimum,\space \frac{3x_1^2x_2}{x_1^3}=\frac{p_1}{p_2}⇒\frac{3x^2}{x_1}=\frac{p_1}{p_2}\\⇒x_2=\frac{p_1}{3p_2}x1"

Putting value of x2ย in x1, we get

"⇒p_1x_1+p_2(\frac{p_1}{3p_2})x1=m\\⇒p_1x_1+\frac{p_1}{3}x1=m\\⇒3p_1x_1+p_1x_1=3m\\⇒4p_1x_1=3m\\⇒x1*=\frac{3m}{4p_1}"


b) Similarly,

"x2^*=\frac{p_1}{3p_2}(\frac{3m}{4p_1})\\x2^*=\frac{m}{4p_2}"

Therefore, the agent's Marshallian demandย "(x1^*,x2^*)= (\frac{3m}{4p_1},\frac{m}{4p_2})"


c) Hence, indirect utility function is derived by putting x1* and x2* in utility function is

"V(m,p1,p2)=(\frac{3m}{4p_1})^3(\frac{m}{4p_2})\\V(m,p1,p2)=\frac{3m^4}{16p_1p_2}"